a machine that fills quart milk is set up to average 32.2 oz per carton, with a standard deviation of 1.2 oz. assuming the distribution is normal find the following:
if five cartons are selected at random, what proportion will have less than 32oz of milk?

Answer by stanbon(75887)

a machine that fills quart milk is set up to average 32.2 oz per carton, with a standard deviation of 1.2 oz. assuming the distribution is normal find the following:
if five cartons are selected at random, what proportion will have less than 32oz of milk?
-----
t(32) = (32-32.2)/(1.2/sqrt(5)) = -0.3726
----------------------
P(x < 32) = P(t < -0.3726 when df = 4) = 0.3642 or 36.42%
========================================================
Cheers,
Stan H.